solve by substitution x=y-2 the second part is 4x+y=12, answer using (___,___)
Answer by Cromlix(4381)

Hi there,
x = y - 2..........(1)
4x + y = 12........(2)
Substitute x = y - 2 from
Eq(1) into Eq(2)
4x + y = 12
4(y - 2)+ y = 12
Multiply out
4y - 8 + y = 12
Collect like terms:
4y + y = 12 + 8
5y = 20
y = 4
Substitute y = 4 into
Eq(1)
x = y - 2
x = 4 - 2)
x = 2
(2,4)
Hope this helps :-)